determine at what temperature aluminum will have the same resistivity as tungsten does at 24 degrees Celsius
Answer #1

The relation between resistivity and temperature is, 5.6x10-s = (2.65x10)[14( 0.00429)(7-24)] 1+(0.00429)(7-24)= 2.11 (0.004